OpenStudy (anonymous):

Convert the polar equation to rectangular form. r=2/1+sin(theta)

OpenStudy (jamesj):

Well, how do you write r in terms of Cartesian coordinates x and y? And then how do you write theta?

OpenStudy (jamesj):

...and sin(theta)?

OpenStudy (anonymous):

I'm not really sure.

OpenStudy (jamesj):

You know that x = r . cos theta y = r . sin theta What are the equations the other way r = .... theta = .... sin(theta) = ... cos(theta) = ...

OpenStudy (jamesj):

r = sqrt(x^2 + y^2), yes? What about theta = .... ?

OpenStudy (jamesj):

Look, we have that x = r . cos theta y = r . sin theta Hence sin theta = ... ?

OpenStudy (anonymous):

y?

OpenStudy (jamesj):

y = r.sin(theta) hence sin(theta) = y/r Yes?

OpenStudy (anonymous):

Yep, got that :)

OpenStudy (jamesj):

So now, substitute these expressions into the original equation \[ r = \frac{2}{1 + \sin\theta} \]

OpenStudy (anonymous):

Alright, I did. I just don't see how I can possibly get to the answer: "x^2 +4y -4"

OpenStudy (jamesj):

Let me do a couple of steps for you r = 2/(1 + sin theta) r = 2/(1 + y/r) r(1 + y/r) = 2 r + y = 2 sqrt(x^2 + y^2) = 2 - y Now, you take it from there

OpenStudy (jamesj):

ugh. Cat just stepped all over keyboard. Writing it out again r = 2/(1 + sin theta) r = 2/(1 + y/r) r ( 1 + y/r) = 2 r + y = 2 r = 2 - y Now you take it from there

OpenStudy (anonymous):

That's the part I always get stuck on though. :( sorry.

OpenStudy (jamesj):

What is r in terms of x and y?

OpenStudy (anonymous):

I end up with 2y + x -2 = 0 and that's not the answer.

OpenStudy (anonymous):

x+y

OpenStudy (jamesj):

Wrong. r is the radial distance of a point (x,y) from the origin. What is the distance of (x,y) from the origin? Use Pythagorus's theorem

OpenStudy (anonymous):

sqrt(x^2+y^2)

OpenStudy (jamesj):

Yes. Hence r = 2 - y is sqrt(x^2 + y^2) = 2 - y Simplify that.

OpenStudy (anonymous):

That's the part I'm stuck on.

OpenStudy (jamesj):

Square roots signs are always pesky. Square both sides to get rid of it.

OpenStudy (anonymous):

Finished the problem. :) Thanks so much!!

OpenStudy (jamesj):

good. Draw a diagram of the figure so you know what this looks like and make sure it jives with your understanding of r and theta.